OpenStudy (anonymous):

I need some help with this question: Which of the following best define 2/(3^3) - Square root of 9 - Cube root of 9 - Cube root of 3 - Square root of 3 I believe the answer is cube root of 3 just because that's what my gut is feeling...

OpenStudy (zzr0ck3r):

\(\dfrac{2}{3^3}\)?

OpenStudy (anonymous):

Yes :)

OpenStudy (zzr0ck3r):

are you sure?

OpenStudy (zzr0ck3r):

\(\dfrac{2}{3^3}=\dfrac{2}{27}\) which is not on the list

OpenStudy (zzr0ck3r):

are you sure its not \(3^{\frac{2}{3}}\)?

OpenStudy (zzr0ck3r):

\(3^{\frac{2}{3}}=\sqrt[3]{3^2}=\sqrt[3]{9}\)

OpenStudy (anonymous):

oooooh - okay I got it now

OpenStudy (anonymous):

Thank you :)